OpenStudy (ksaimouli):

solve each system x-9y=25 6x-5y=3

OpenStudy (angela210793):

x-9y=25(*6) 6x-54y= 150 (-) both eq -49y=147 6x-5y=3 6x-5y=3 y=-147/49=3 replace y in the 1st or 2nd eq 6x-15=3 x=18/6=3

OpenStudy (anonymous):

x=25+9y now substitute x in 2nd equation 6(25+9y)-5y=3 150+54y-5y=3 49y=3-150 y=147/49

OpenStudy (anonymous):

y=-3 now substitute y value in one of the equation x=25+9(-3) x=25-27 x=-2

OpenStudy (karatechopper):

ajmal! mere dost!